In need of help!!
Simplify.[2 × (10 + 5)] - 5

Found 3 solutions by rfer, Math-Genius, stanbon:
Answer by rfer(16322)

(2x(10+5))-5
2x15-5
30-5=25
Follow the order of operations.

Answer by Math-Genius(4)

First, add what is in the parenthesis.
(10+5)= 15
Second, fill in 15 in the equation where (10+5) was.
[2*15]-5
Third, multiply 2*15.
2*15=30
Fourth, fill in 30 where [2*15} was.
30-5
Last, subtract 5 from 30.
30-5=25
Then you have your answer.

Answer by stanbon(75887)

Simplify.[2 × (10 + 5)] - 5
-------
= [2*(15)]-5
-------
= [30]-5
-------
= 25
==========
Cheers,
Stan H.